Determine the following for a 15 g sample of carbon monoxide gas:

a. Pressure in mmHg, at 35 degrees Celcius in a 50 L container

b. Volume in L at 127 degrees Celcius and 3.0 atm

c. Temperature in degree Celcius, at 5.0 atm pressure and 5.0 L


1
Expert's answer
2021-11-29T04:37:45-0500

Molar mass of carbon monoxide=28.01

15/28.01=0.54mol

PV=nRT

P×50=0.54×8.314×308

50P=1371.32

P=27.43